James bought 80 postage stamps for which he paid $3.50. Some were 3-cents stamps and some were 5-cents stamps. How many of each kind did he buy?
Answer by checkley77(12844)

x+y=80 or x=80-y
.03x+.05y=3.50
.03(80-y)+.05y=3.50
2.40-.03y+.05y=3.50
.02y=3.50-2.40
.02y=1.10
y=1.10/.02
y=55 $.05 stamps.
x=80-55=25=$.03 stamps.
Proof:
.03*25+.05*55=3.50
.75+2.75=3.50
3.50=3.50